Solve each system by substitution:
5x+y=10
3x-3y=6

Answer by checkley77(12844)

5x+y=10
y=-5x+10
3x-3y=6
3x-3(-5x+10)=6
3x+15x-30=6
18x=6+30
18x=36
x=36/18
x=2 ans.
y=-5*2+10
y=-10+10
y=0 ans.
Proof:
3*2-3*0=6
6=6